OpenStudy (anonymous):

the population of a bacteria grows continuously at 15% per hours suppose there are 134 bacteria at time t=0 how long will the bacteria doubles

OpenStudy (anonymous):

solve \[(1.15)^t=2\] for \(t\)

OpenStudy (anonymous):

this takes only one step and a calculator \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] in your case \(b=1.15,x=t, A=2\)